User Tools

Site Tools


operations:agenda2016_12_22

Chair/Presenting: Tiege

In Attendance: Tiege, Ellen, Nick, Jim, Lucia, Liz, Jamie, Ross, Lucas, Bryn

Apologies: Jesse, Jonny, Arwin

Absent:

___

Agenda:

General Issues

  • Fresh roster out for January, reminder that some shifts were swapped around a day or so after the roster was posted so double check yo shifts.
  • LBA session is starting on the 18th of January, we will be covering LBA while rostered on for AuScope. If people haven't had to cover any LBA yet, might need to do a refresher on common problem troubleshooting etc. for Ho and Cd.
    • Organise a refresher session prior to the observations. TBA.
  • Jim's post glitch observations, mostly wrapped up now?
    • No more expected issues, vela observations are lower priority now.
  • Christmas drinks?
    • METZ 12:30pm! Leave at 12:15pm.
  • New log monitor bug testing.
    • People should give it a go as the default, and report any issues. Can fall back on the old one if needed.
    • Cd will be getting added to the log monitor and gaining ERC support. IMPORTANT: DETERMINE COLOUR FOR CD. Purple decided.
  • Brett and co. doing some work on the Y bearing (Ho 26m) from the 6th of January for two weeks, hopefully will be finished earlier than this. Chance RD1701 will be impacted by this.
  • Can restart Dave's delay monitoring script in order to only include data from the current experiment.
  • IF YOU WANT TO DO THE ROSTERING SPEAK UP, Talk to Jim/Ellen.

Issues from Handover notes

Does not appear to be a large number of recurring problems. Haven't been here for the last two weeks.

  • Hobart 12m getting stuck drives/not recovering from wind stows a few times the last two weeks. (R1770, AUA015)
  • Some reports of dodgy internet connection to Yg that resetting the VPN does not fix.
    • power cycling router tends to work.
/home/www/auscope/opswiki/data/pages/operations/agenda2016_12_22.txt · Last modified: 2016/12/22 02:26 by Tiege McCarthy